A carnot engine working between 27c and 127c takes up 800j of heat drom the reservoir in one cycle.what is the work done by engine?

Grade:12

2 Answers

Arun
6 years ago
Dear Dheeraj
 
As we know
Efficiency (n) = 1 - T2/T1
= 1 -300/400
= ¼
Now
n = W/Q1
Hence
Work done = efficiency (n) * heat absorbed(Q1)
= ¼ * 800
= 200 Joules
 
Regards
Arun (askIITians forum expert)
SHAYAN NATH
4 years ago
As we know efficiency (n)= w/Q
 So we write n= 1-T2/T1= 1-300/400
                                          = 1-3/4 = 0.25=25%
    We know, w=Q×n= 800×0.25=200(Answer)